# Tractable Equilibrium Computation in Markov Games through Risk Aversion

### v1

@@ 1 Introduction
Machine learning algorithms are increasingly being deployed in dynamic environments in which they interact with other agents like people or other algorithms. Often, these agents have their own goals which may not be aligned with those of the algorithm—making the interactions strategic. These situations are naturally modeled as games between rational agents and their prevalence in a wide range of application areas has driven a surge in research interest in learning in games [10] and multi-agent reinforcement learning [78] in recent years. Indeed, real-world applications of these problems range from the training of large language models [47] to the decentralized control of the smart grid [46], autonomous driving [39], and financial trading [74] among many others.
Due to the game-theoretic structure of the problem, many of these problems can be cast as a problem of equilibrium computation under varying information structures, where the equilibrium represents a stable outcome for rational agents. The most common solution concept is that of a Nash equilibrium (NE) [48]: a solution under which no rational agent has an incentive to unilaterally seek to improve their outcome. Despite its popularity as a solutation concept, computing a NE outside of highly structured games is known to be computationally intractable [11] even for two-player matrix games.
While relaxations of NE like (coarse) correlated equilibria (CCE) are known to be more tractable to compute, they also have their limitations. Indeed while CCE can be computed through the use of no-regret learning algorithms [10], the set of CCE can be large (introducing an additional problem of equilibrium selection) and may include strictly dominated strategies [71] which means that cannot necessarily be rationalized by individual agents [14]. Furthermore, a refinement of CCE—stationary CCE— are also known to be PPAD-hard to compute in dynamic general-sum games [13].
Beyond these hardness results, solution concepts like NE and CCE also fail to be predictive of what strategies people play in games [42, 16], with people being observed to be imperfect optimizers [26, 8] and risk-averse [27] when confronted with game theoretic scenarios. This aligns with celebrated work in behavioral economics and mathematical psychology which has repeatedly shown that dominant features of human decision-making are risk-aversion [37, 70] and a failure to perfectly optimize [41].
The first observation is often attributed to an agent’s bounded rationality which posits that individuals are naturally prone to making mistakes and often fail to be perfectly optimal [41]. This is often captured in games through the idea of a quantal response equilibrium (QRE) [42].
@@
These results suggest that designing MARL agents with realistic features of human decision-making can allow for a computationally tractable equilibrium concept in all finite action games and finite horizon Markov games. To emphasize the practical relevance of this theoretical result, we show how that the regime for which RQE are computationally tractable captures real-world data from behavioral economics on people’s observed pattern of play in $13$ different games. This is illustrated in Fig. 1, where the blue region represents the set of RQE that are computationally tractable which is a function of agents’ degree of risk-aversion ($\tau_{1}/\tau_{2}$), and level of bounded rationality ($\epsilon_{1}/\epsilon_{2}$).
@@
To begin, in Section 2 we focus on multi-player general-sum matrix games. We first introduce risk-aversion into matrix games by allowing agents to be risk-averse to the randomness introduced into the game by their opponents. To incorporate risk-aversion we make use of a general class of convex risk measures. We show that for any convex risk measure, a risk-averse Nash equilibrium must exist. We then introduce bounded rationality by assuming that agents optimize over quantal responses instead of the entire probability simplex. In Section 3 we show that computing the QRE of risk-averse matrix games—i,e., the RQE described above—can be done in polynomial time as long as the level of risk-aversion and class of quantal response functions satisfies certain conditions Importantly, these conditions are independent of the underlying game and only depend on the class of risk metrics and quantal responses under consideration. Thus, RQE not only incorporates features of human decision-making but are also more amenable to computation than QRE or NE in matrix games.
In Section 4, we extend the study to multi-player general-sum Markov games. In this setting we consider the case in which players are risk-averse not only to the randomness introduced by their opponents, but also to the randomness in the environment. We again show that RQE can be computed efficiently when the environment is known and one has full information over the game. In addition, we consider the multi-agent reinforcement learning (MARL) setting in which the environment is unknown and one needs to learn by sampling. We provide a first algorithm and a finite-sample guarantee for learning an approximate RQE in such MARL problems. Finally, we conduct experiments of the dynamic games in gridworlds to illustrate the effect of risk-aversion in sequential games and validate our findings.
@@ 1.1 Related Works
Before presenting our results, put our work into context, and discuss related works.
Computation tractability of NE in games.
Computing NE generally is known to be PPAD-hard [11] even for two-player general-sum matrix games. Many long lines of work study the approximation of NE equipped with no-regret learners in some structured games. In structured games, one typically makes assumptions on the properties of the utilities in the games which make NE computation feasible. We summarize representative works here that propose NE computation in structured games. Concave and Monotone Games [22, 33]: [19] show there exist no-external regret methods to achieve an NE in socially concave games (a sub-class of concave games). [30] generalize to monotone (and smooth) games equipped with no-regret methods to achieve a NE. Coarse Correlated Equilibria (CCE) [72, 3]: [7, 38] show that the CCE implies NE in zero-sum polymatrix normal-form matrix games and Markov games respectively. The containment of the set of NE in the set of CCE holds trivially by definition. Quantal Response Equilibria (QRE) [42, 43, 45]: When the payoffs are perturbed by Gumbel-distributed noise, the NE of such games are logit QRE, thus extending the solution concept. More general noise and their associated QRE are realized by the games with payoffs regularized by the quantal response functions (Definitions 4 and 2). [9, 69] provide no-regret methods to attain logit QRE and KL-regularized QRE respectively. Smooth Payoffs and Equilibria [59, 4, 12]: [4] showed no-regret methods on smoothified (w.r.t the payoffs) games that approximate NE. [12] instead proposed strong smooth equilibria of a normal-form game that collapses to the set of Nash equilibria in the polyhedral concave game which depends on the smoothness parameter. While such computationally tractable algorithms exist to approximate NE, both NE and CCE fail to be predictive of what strategies people play in games [42, 16]. Thus, we first introduce the risk-adjusted Nash equilibria (RNE) solution concept (c.f. Definition 3) for general-sum matrix games to capture real-world decision-making strategies. We also provide computationally tractable methods to get to these equilibria. To emphasize, different from the previously mentioned works, we provide a principled optimization approach (c.f. Section 3) that is an RNE-computable for general-sum matrix games.
Risk-averse single-agent Reinforcement Learning:
[64] proposed the framework of risk-sensitive RL, whose goal is learning risk-averse policies, by defining a ‘valuation function’ that satisfies two important properties: monotonicity and translation invariance. These risk-averse policies are aversions to the risks originating from environmental uncertainties. Many lines of work consider specialized valuation functions: The vanilla class of utility-based shortfalls [21] are actually valuation functions. [34, 20] considers the exponential utility function, [15, 2, 55] considers various value-at-risk risk functions, and in the context of robust reinforcement learning that alleviate environmental uncertainties, [35, 49, 80, 50, 65, 52, 75, 66, 53, 51, 57] considers distributional robust optimization-based functions which are valuation function. [31] considers both aleatory (environmental) uncertainties and epistemic (model estimation) uncertainties using entropic value-at-risk valuation function. Different from these works, we consider the risk-aversion solution concept (Definition 7) to address both multi-agents and aleatory uncertainties (c.f. Section 4).
Risk-averse/robust multi-agent Reinforcement Learning:
Risk-aversion [27, 61, 54] has been showcased as a promising strategy choice that aligns with how people play games. Preliminary works such as [76, 73] and [24, 17, 68] proposed concepts of risk-averse equilibria solution in games and Markov games, respectively. They typically rely on structural assumptions, like competition [76] or monotonicity [73] structures, directly based on risk-averse adjusted games with risk-averse utilities. We show such structures may not exist for risk-averse adjusted games in Example 1. Many risk-averse multi-agent deep reinforcement learning works [23, 79, 56, 63, 68] showcase promising empirical guarantees of heuristic risk-averse objectives, highlighting risk-averse policy learning. Preliminary works on robust multi-agent reinforcement learning [77, 32, 67] focus on addressing environmental uncertainty issues, assuming the existence of NE in robust Markov games, through their practical sampling-based algorithms. More recently, [44] additionally considers the multi-agent and environmental uncertainties, and provides a collapse to the zero-sum NE by making structural assumptions on the regularizers. In contrast to these above-mentioned works, we provide a principled approach to solving risk-averse equilibria with a computationally tractable and sample-efficient algorithm. Perhaps more closely related to equilibria concepts (Definition 3) of our work, [36, 18] consider KL-regularized NE concepts with the usual competitive utility objective (c.f. Eq. 1). In particular, the KL-regularization helps bias the player policy towards a human-imitation learned anchor policy, hoping for bounded rational behaviors. [36] shows, in the case of a two-player zero-sum game, that such KL-regularized NE collapses to approximate (up to constant based on the KL regularized parameter) NE of the original utility objective (c.f. Eq. 1). In more generality, our work makes a shrewd observation of combining both risk-aversion and bounded rationality concepts to achieve computational tractability (c.f. Section 3) for an $n$-player general-sum game.
@@ 2.1 Risk-Aversion in Matrix Games
In the case of matrix games, we consider the case where agents are risk averse with respect to the mixing or randomness introduced into the problem by their opponents. In generalizing this to stochastic games in future sections we will also consider the case where agents are risk averse with respect to the underlying dynamics in the Markov game. This is a common approach taken in the literature on risk-sensitive and robust decision-making [64]. We refer to related works for more discussion. Thus, we transform the player’s utilities in (1) into costs $f_{i}$ which take the form
$\begin{split}f_{i}(\pi_{i},\pi_{-i})=\mathbb{E}_{\pi_{i}}[\rho_{i,\pi_{-i}}(R_{i}(a))]=\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i})\rho_{i,\pi_{-i}}(R_{i}(a_{i},a_{-i}))\end{split}$ (3)
where $\rho_{i,\pi_{-i}}$ is used to denote the potentially different risk preference of agent $i$ which depends on the product distribution of opponents strategies $\pi_{-i}$. Note that if $\rho_{i}(X)=\mathbb{E}[-X]$ for all players $i=1,...,n$ (which satisfies the requirements in Definition 2) then the new formulation reduces to the original expected utility objective. Thus, this formulation can be seen as a strict generalization of the classic setup described in the previous section.
While—at first glance— this modified game looks significantly more complex than the previous expected utility maximization setup (cf. (1)), we can rely on a particularly powerful property of convex measures of risk to simplify and expose some structure in this class of problems. In particular, we make use of the dual representation theorem for convex risk measures.
@@
Given this setup, we can now write the risk-adjusted game in the following form:
@@
where $\mathcal{P}_{-i}=\mathcal{P}/\Delta_{A_{i}}\subset\mathbb{R}^{A_{-i}}$, $A_{-i}=\sum_{j\neq i}A_{j}$, and $R_{i}\in\mathbb{R}^{A_{i}\times A_{-i}}$ is player $i$’s payoff matrix.
We note that we differentiate the penalty functions $D_{i}$ to allow for the different agents to have different risk preferences. In this form, one can see that in a risk-adjusted game, the players imagine that an intermediate adversary who seeks to maximize their cost within a small ball in some metric of the opponents realized strategies. Thus, agents in this class of games introduce a certain amount of worst-case thinking into their strategies. Note that from the point of view of player $i$ the game has simplified from a setting with $n-1$ opponents to a game with $1$ opponent who is penalized from taking actions that deviate too far from the product distribution of the original $n-1$ opponents’ strategies.
Since the Nash equilibrium of the risk-adjusted game can be qualitatively different from that of the original game, we define a risk-adjusted Nash equilibrium (RNE) as follows.
@@
A risk-adjusted Nash equilibrium of is joint strategy $\pi^{*}\in\mathcal{P}$ such that no player has any incentive to unilaterally deviate in the risk adjusted game—i.e., for all $i=1,...,n$
@@
Note that since players would like to minimize risk, the direction of the inequality has changed. The convexity and continuity of the penalty function guarantees that the risk-adjusted game admits at least one RNE.
@@
There exists a RNE for the risk-adjusted game (4).
@@
To begin, we show that $f_{i}(\pi_{i},\pi_{-i})$ is convex in $\pi_{i}$ for all $\pi_{i}\in\mathcal{P}_{-i}$. This follows from the fact that $D_{i}(\cdot,\cdot)$ is assumed to be convex and continuous in its first argument. Invoking Danskin’s theorem guarantees us that $f_{i}$ is convex in $\pi_{i}$ for all fixed values of $\pi_{-i}$. Since the probability simplex is compact and convex, the game satisfies the conditions of a convex game [58] and thus a Nash equilibrium must exist. ∎
Despite the fact that risk preferences already help convexify player’s objectives, the addition of risk can serve to weaken existing structures in the player’s cost function. In particular, we show that even in two-player zero-sum games where $R_{1}=R=-R_{2}^{T}$, the risk-adjusted game loses any zero-sum structure and may cease to even be strictly competitive in the sense that one player’s gain is the other’s loss.
@@
Consider a 2-player zero-sum game where $R_{1}=R=-R_{2}^{T}$ where players are risk averse in the entropic risk metric with different degrees of risk aversion $\tau_{1}$ and $\tau_{2}$. Under these conditions, the players loss functions take the following form:
@@
Even instantiating $R=\mathbb{I}_{2}$, $\tau_{1}\geq 10,\tau_{2}>0$, both $f_{1}(\pi_{1},\pi_{2})>f_{1}(\pi_{1}^{\prime},\pi_{2})$ and $f_{2}(\pi_{1},\pi_{2})>f_{2}(\pi_{1}^{\prime},\pi_{2})$ holds for the regions $\pi_{1},\pi_{2}\in\Delta_{2}$ satisfying $\pi_{2}(1)\in(0.1,0.5)$ and $0.75-\pi_{1}(1)>\pi^{\prime}_{1}(1)>\pi_{1}(1)$, which implies that the game is not a strictly competitive game.
@@
Finally, note that the additional convexity induced by the introduction of risk aversion guarantee is not enough to ensure the computational tractability of the Nash equilibrium. This is not surprising given the fact that computing NE in general convex games can be intractable and that further structural conditions are required for computational tractability.

### v2

@@ 1 Introduction
Machine learning algorithms are increasingly being deployed in dynamic environments in which they interact with other agents like people or other algorithms. Often, these agents have their own goals which may not be aligned with those of the algorithm—making the interactions strategic. Such situations are naturally modeled as games between rational agents and their prevalence in a wide range of application areas has driven a surge in research interest in learning in games [13] and multi-agent reinforcement learning [76] in recent years. Indeed, real-world applications of these problems range from the training of large language models [50] to the decentralized control of the smart grid [48], autonomous driving [39], and financial trading [72] among many others.
When viewed through the lens of game-theory, many of these problems can be cast as problems of equilibrium computation under varying information structures, where the equilibrium represents a stable outcome for rational agents. The most common equilibrium concept is that of a Nash equilibrium (NE) [51]: a solution under which no rational agent has an incentive to unilaterally seek to improve their outcome. Despite its popularity as a solution concept, computing a NE outside of highly structured games is known to be computationally intractable [14] even for two-player matrix games.
While relaxations of NE like (coarse) correlated equilibria (CCE) are known to be more tractable to compute, they also have their limitations. Indeed while CCE can be computed through the use of no-regret learning algorithms [13], the set of CCE can be large (introducing an additional problem of equilibrium selection) and may include strictly dominated strategies [70] which means that they cannot necessarily be rationalized by individual agents [17]. Furthermore, a refinement of CCE—stationary CCE— are also known to be PPAD-hard to compute in dynamic general-sum games [16].
Beyond these hardness results, solution concepts like NE and CCE also fail to be predictive of what strategies people play in games [44, 18], with people being observed to be imperfect optimizers [27, 11] and risk-averse [28] when confronted with game theoretic scenarios. This aligns with celebrated work in behavioral economics and mathematical psychology which has repeatedly shown that dominant features of human decision-making are a failure to perfectly optimize [42] and risk-aversion [37, 69].
The first observation is often referred to as bounded rationality which posits that individuals are naturally prone to making mistakes and often fail to be perfectly optimal [42]. This is often captured in games through the idea of a quantal response equilibrium (QRE) [44].
@@
These results suggest that designing multi-agent reinforcement learning (MARL) agents with realistic features of human decision-making can allow for a computationally tractable equilibrium concept in all finite action games and finite horizon Markov games. To emphasize the practical relevance of this theoretical result, we show how that the regime for which RQE are computationally tractable captures real-world data from behavioral economics on people’s observed pattern of play in $13$ different games. This is illustrated in Fig. 1, where the blue region represents the set of RQE that are computationally tractable which is a function of agents’ degree of risk-aversion ($\tau_{1}/\tau_{2}$), and level of bounded rationality ($\epsilon_{1}/\epsilon_{2}$).
@@
To begin, in Section 2 we focus on multi-player general-sum matrix games. We first introduce risk-aversion into matrix games by allowing agents to be risk-averse to the randomness introduced into the game by their opponents. To incorporate risk-aversion we make use of a general class of convex risk measures and present two forms of risk aversion: aggregate risk-aversion and action-dependent risk aversion—with the latter yielding less conservative solutions at the cost of a more complex formulation and analysis. Given this setup, we show that risk-averse Nash equilibria must exist in both formulations in all games and for any convex measures of risk.
We then introduce bounded rationality into the games by assuming that agents optimize over quantal responses instead of the entire probability simplex. In Section 3 we show that computing the QRE of both types of risk-averse matrix games—i,e., the RQE described above—can be done in polynomial time as long as the level of risk-aversion and class of quantal response functions satisfies certain conditions. Importantly, these conditions are independent of the underlying game and only depend on the class of risk metrics and quantal responses under consideration. Thus, RQE not only incorporates features of human decision-making but are also more amenable to computation than QRE or NE in matrix games.
In Section 4, we extend the study to multi-player general-sum Markov games. In this setting we consider the case in which players are risk-averse not only to the randomness introduced by their opponents, but also to the randomness in the environment. We again show that RQE can be computed efficiently when the environment is known and one has full information over the game. In addition, we consider the multi-agent reinforcement learning setting in which the environment is unknown and one needs to learn by sampling. We provide a first algorithm and a finite-sample guarantee for learning an approximate RQE in such MARL problems. Finally, we conduct experiments of the dynamic games in gridworlds to illustrate the effect of risk-aversion in sequential games and validate our findings.
@@ 1.1 Related Works
Before presenting our results, we put our work into context, and discuss related works here.
Computational tractability of game theoretic solution concepts.
This work proposes a new solution concept for game theoretic settings that is computationally tractable, yet retains many of the desirable properties of classical equilibrium concepts. This general question emerged from the finding that computing a Nash equilibrium—perhaps the most natural solution concept for a game between rational self-interested agents—is PPAD-hard [14], even for two-player general-sum matrix games. Despite this negative result, a large amount of subsequent work has focused on understanding the classes of games in which one can compute, approximate, or learn Nash equilibria efficiently. This is often done by assuming additional structure on the players’ utilities and their relationships to one another, with large classes of games being zero-sum or competitive games, zero-sum polymatrix games [10, 38], monotone games [31], smooth games [57], or socially concave games [21].
In games without such structure however, the natural targets for computation and learning became correlated [49] and coarse correlated equilibria  [4, 5](CE and CCE respectively), both of which can be shown to emerge as the endpoint of no-regret learning and are thus considered to be computationally tractable targets for the design of learning algorithms. Despite this desirable property, the two concepts have significant drawbacks. Indeed both CE and CCE require some form of coordination between players to implement, introduce a highly nontrivial equilibrium selection problem [13], and may have support on dominated strategies [70]. Furthermore, in the dynamic game context of Markov games, stationary CE and CCE are also computationally intractable to compute [16].
More recently, a new equilibrium concept—a smoothed Nash equilibrium— has been proposed as an alternative to these other equilibrium concepts [15] and motivated by similar considerations of individual and independent rationalizability and computational tractability. By applying ideas from smoothed analysis to the problem of computing Nash equilibria the authors show that one can efficiently find approximate classes of smoothed Nash equilibria—though to the best of our knowledge this cannot be done in a decentralized way.
Our approach is orthogonal and is rooted in giving MARL agents a foundation rooted in behavioral economics by imbuing them with a realistic feature of human decision-making: risk-aversion. The question of computational tractability of risk-averse Nash equilibria has been analyzed in [22]. The work shows that if agents are risk-averse with respect to all the randomness in the game (including their own) then a risk-averse Nash equilibrium may not even exist in mixed strategies, and even understanding if such equilibria exist can be NP-complete. Our formulation overcomes this by incorporating risk-aversion in a different way. Indeed, we show that when agents are risk-averse only to the randomness introduced to their opponents (and the environment) then the risk-averse Nash equilibria will always exist. We note that such formulations of risk-aversion are common in the literature on risk-sensitive control [62, 8] and risk-sensitive reinforcement learning [63] where agents are implicitly presumed to be risk-averse only to the randomness that is outside their control (i.e., the environment). Furthermore we show that introducing bounded rationality into the game allows a class of risk-averse quantal response equilibria (RQE) to be computationally tractable in all finite action and finite-horizon Markov games.
Predictive power of equilibrium concepts.
Another driving force in moving beyond the Nash and correlated equilibrium concepts stems from their lack of predictive power in experimental settings (see e.g.,  [9, 52, 43, 18, 44]). To address this, a line of work originating in economics seeks to understand the natural solution concepts in game where players have behaviorally plausible restrictions to their strategy spaces, and to study whether such equilibria were better predictors of human play than Nash or (coarse) correlated equilibria [28, 29, 33]. The most common restriction is that players have bounded rationality,—i.e., they may fail to perfectly optimize—a model with roots in mathematical psychology [42]. Under this restriction, a natural equilibrium concept that emerged was that of a quantal response equilibrium (QRE) which induces bounded rationality by either assuming that the players are rational in a stochastically perturbed version of the game or equivalently that they optimize a regularized version of their utility [44, 45, 47]. Beyond their use as a better model for human decision-making in games, QRE have also increasingly been adopted as a solution concept in multi-agent reinforcement learning and learning in games [68, 47, 12, 41, 20, 35] due to their links with KL and entropy regularized reinforcement learning. Despite these developments QRE are not computable in all games. Indeed the class of QRE or equivalently the level of bounded rationality needed for computational tractability depends on the underlying game structure which may not be known a priori. In contrast we show that the addition of risk aversion allows for the same class of quantal response equilibria to be computationally tractable to compute in all finite action games and finite-horizon Markov games. Furthermore we show that this class of risk-averse QRE is nontrivial and can capture human data better than risk-neutral QRE—a finding which is in line with findings in behaviorial economics [28, 29].
Risk-averse and robust multi-agent reinforcement learning.
Our work builds on and provides an additional justification for risk-sensitive (multi-agent) reinforcement learning. This line of work has roots going back to seminal work by Jacobson on risk-sensitive control [36], and more recently in risk-sensitive reinforcement learning [63]. In these works, the aim is to find a controller or policy for a system that accounts for stochasticity or uncertainty in the environment or system in a more nuanced way than risk-neutral approaches like optimal control or reinforcement learning [8]. Due to classic duality results (see e.g.,  [77]) this line of work is closely related to the literature on robust control and distributionally robust reinforcement learning [34, 53, 64, 73] which seeks to find solutions that are robust to worst case environmental disturbances.
Our work rigorously extends these formulations to the multi-agent regime though it is not the first to consider risk-aversion in MARL. Indeed, risk-sensitive MARL has been the focus of several recent works (e.g., [74, 71, 25, 67]). Several provide rigorous definitions of risk-averse equilibria and some guarantees on their computation by assuming structure on the risk-adjusted game. Oftentimes this is done by assuming that the risk-adjusted game is itself zero-sum [74], monotone [71], or that it satisfies other strong conditions [25]. Other works are more empirical in nature [24, 19, 78, 55, 61, 67], showing the promise of risk-averse algorithms for MARL.
One last closely related line of work is the emerging literature on robust multi-agent reinforcement learning [75, 32, 66, 7]. Once again due to duality arguments, these works can be seen as tackling a similar problem to the risk-averse MARL problem. The focus of these previous works, however, is on robustness in the face of only environmental uncertainties (and not opponent strategies), and questions of existence and computational tractability are either assumed away or the focus is on extensions of correlated equilibrium concepts. A recent related work in this literature analyzed the computational tractability of robust Nash equilibria in Markov games, but only provided strong guarantees on the zero-sum regime, showing that computing such equilibria in general is PPAD-hard [46].
To the best of our knowledge, no previous work in either of these literatures highlights the broad benefits afforded by risk-aversion in MARL in terms of computational tractability of equilibria. In our work we show that risk-aversion (and by extension distributional robustness), when combined with bounded rationality yields a computationally tractable class of individually rationalizable equilibria in all finite-horizon $n$-player Markov games. Furthermore we show that these equilibria can be computed using no-regret learning algorithms.
@@ 2.1 Risk-Aversion in Matrix Games
In the case of matrix games, we consider the case where agents are risk averse with respect to the mixing or randomness introduced into the game by their opponents. In generalizing this to stochastic games in future sections we will also consider the case where agents are risk averse with respect to the underlying dynamics in the Markov game. Crucially, we assume that players are not risk averse to their own randomness. We note that is a common approach taken in the literature on risk-sensitive and robust decision-making [63] and it is necessary since if agents are risk-averse to their own randomness then an equilibrium may cease to exist [22]. We refer to related works for more discussion.
We consider two generalizations of the original game that differ in how risk is incorporated into the problem: aggregate risk aversion and action-dependent risk aversion. The first generalization allows for a comparatively simpler analysis but at the cost of a more conservative solution while the second is less conservative but requires a more complex analysis. Due to monotonicity and linearity of expectation both of these formulations can be seen as generalizations of risk-sensitive decision-making investigated in the literature on risk-sensitive control [8] and risk-sensitive RL [63]. To see this reduction, one can simply take the other agents to be part of an unknown environment.
Aggregate Risk Aversion:
To define a game with aggregate risk aversion, we transform the player’s utilities in (1) into costs $f_{i}$ which take the form
$\begin{split}f_{i}(\pi_{i},\pi_{-i})=\rho_{i,\pi_{-i}}(\mathbb{E}_{\pi_{i}}[R_{i}(a)])=\rho_{i,\pi_{-i}}\left(\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i})R_{i}(a_{i},a_{-i})\right)\end{split}$ (3)
where $\rho_{i,\pi_{-i}}$ is used to denote the potentially different risk preference of agent $i$ which depends on the product distribution of opponents strategies $\pi_{-i}$.
Action-dependent Risk Aversion:
To define a game with aggregate risk aversion, we transform the player’s utilities in (1) into costs $f_{i}$ which take the form
$\begin{split}f_{i}(\pi_{i},\pi_{-i})=\mathbb{E}_{\pi_{i}}\left[\rho_{i,\pi_{-i}}(R_{i}(a))\right]=\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i})\rho_{i,\pi_{-i}}\left(R_{i}(a_{i},a_{-i})\right)\end{split}$ (4)
where again $\rho_{i,\pi_{-i}}$ is used to denote the potentially different risk preference of agent $i$ which depends on the product distribution of opponents strategies $\pi_{-i}$.
We remark that in both of these formulations, if $\rho_{i}(X)=\mathbb{E}[-X]$ for all players $i=1,...,n$ (which satisfies the requirements in Definition 2) then the new formulation reduces to the original expected utility objective. Thus, both can be seen as generalizations of the classic setup described in the previous section.
While—at first glance— the modified games look significantly more complex than the previous expected utility maximization setup (cf. (1)), we can rely on a particularly powerful property of convex measures of risk to simplify and expose some structure in this class of problems. In particular, we make use of the dual representation theorem for convex risk measures.
@@
Given this setup, we can now write the aggregate risk-adjusted game in the following form:
@@
where $\mathcal{P}_{-i}=\mathcal{P}/\Delta_{A_{i}}\subset\mathbb{R}^{A_{-i}}$, $A_{-i}=\sum_{j\neq i}A_{j}$, and $R_{i}\in\mathbb{R}^{A_{i}\times A_{-i}}$ is player $i$’s payoff matrix. Similarly, the action-dependent risk-adjusted game takes the form:
$\begin{split}f_{i}(\pi_{i},\pi_{-i})=\sum_{j\in\mathcal{A}_{i}}\pi_{i}(j)\left(\sup_{p_{i,j}\in\mathcal{P}_{-i}}-\langle R_{i,j},p_{i,j}\rangle-D_{i}(p_{i,j},\pi_{-i})\right),\\ \end{split}$ (6)
where $R_{i,j}$ corresponds to the $jth$ row of $R_{i}$.
We note that we differentiate the penalty functions $D_{i}$ to allow for the different agents to have different risk preferences. In this form, one can see that in a risk-adjusted game, the players imagine that intermediate adversaries seek to maximize their cost within a small ball in some metric of the opponents realized strategies. Thus, agents in risk-averse games introduce a certain amount of worst-case thinking into their strategies.
In the aggregate case, from the point of view of player $i$ the game has simplified from a setting with $n-1$ opponents to a game with $1$ opponent who is penalized from taking actions that deviate too far from the product distribution of the original $n-1$ opponents’ strategies. In the action-dependent case, the game from player $i$’s point of view goes from a $n$-player game to a $A_{i}$-player game who are also similarly constrained.
Since the Nash equilibrium of the risk-adjusted game can be qualitatively different from that of the original game, we define a risk-averse Nash equilibrium (RNE) as follows. For ease of exposition we do not differentiate between aggregate and action-dependent RNE since similar results hold for either formulation.
@@
A risk-averse Nash equilibrium of is joint strategy $\pi^{*}\in\mathcal{P}$ such that no player has any incentive to unilaterally deviate in the risk adjusted game—i.e., for all $i=1,...,n$
@@
Note that since players would like to minimize risk, the direction of the inequality has changed. The convexity and continuity of the penalty function guarantees that the risk-adjusted games admit at least one RNE.
@@
There always exists a RNE for all aggregate and action-dependent risk-averse games presented in (5) and (6) respectively.
@@
To begin, we show that $f_{i}(\pi_{i},\pi_{-i})$ is convex in $\pi_{i}$ for all $\pi_{i}\in\mathcal{P}_{-i}$ in (5). This follows from the fact that $D_{i}(\cdot,\cdot)$ is assumed to be convex and continuous in its first argument. Invoking Danskin’s theorem guarantees us that $f_{i}$ is convex in $\pi_{i}$ for all fixed values of $\pi_{-i}$. Since the probability simplex is compact and convex, the game satisfies the conditions of a convex game [56] and thus a Nash equilibrium must exist. The action-dependent risk aversion regime follows from the linearity of $f_{i}$ in $\pi_{i}$ in (6) and a similar invocation of the existence of Nash in convex games [56]. ∎
The fact that these Nash equilibria exist is a consequence of the fact that players are not risk averse to the randomness that they introduce into the game by playing a mixed strategy. Under the stronger assumption that they are risk-averse to all randomness, previous work [22] shows that RNE may not exist. Thus, our specific formulation of risk is crucial in that it will always yield a RNE.
Despite the fact that risk preferences already help convexify player’s objectives, the addition of risk can serve to weaken existing structures in the player’s cost function. To illustrate this, we show that even in two-player zero-sum games where $R_{1}=R=-R_{2}^{T}$, the risk-adjusted game loses any zero-sum structure and may cease to even be strictly competitive in the sense that one player’s gain is the other’s loss.
@@
Consider a 2-player zero-sum game where $R_{1}=R=-R_{2}^{T}$ where players have aggregate risk aversion in the entropic risk metric with different degrees of risk aversion $\tau_{1}$ and $\tau_{2}$. Under these conditions, the players loss functions take the following form:
@@
Even instantiating $R=\mathbb{I}_{2}$, $\tau_{1}=10$, and for any $\tau_{2}>0$, both $f_{1}(\pi_{1},\pi_{2})>f_{1}(\pi_{1}^{\prime},\pi_{2})$ and $f_{2}(\pi_{1},\pi_{2})>f_{2}(\pi_{1}^{\prime},\pi_{2})$ holds for the regions $\pi_{1},\pi_{2}\in\Delta_{2}$ satisfying $\pi_{2}(1)\in(0.1,0.5)$ and $0.75-\pi_{1}(1)>\pi^{\prime}_{1}(1)>\pi_{1}(1)$, which implies that the game is not a strictly competitive game.
@@
Finally, note that the additional convexity induced by the introduction of risk aversion guarantee is not enough to ensure the computational tractability of the Nash equilibrium (see e.g., [46]). This is not surprising given the fact that computing NE in general convex games can be intractable and that further structural conditions are required for computational tractability.